OpenStudy (anonymous):

Hi I just signed up.. pulling my hair out (since yesterday morning - sad i know) trying to figure out this linear equation; An electricity supplier has two tariff rates for charging, and a consumer is free to choose whichever rate is desired; RATE A; 10c/unit for lighting and 2c/unit for power RATE B: a fixed charge of $10 per quarter, plus 3c/unit for either lighting or power. ok, question is... If a consumer uses 1000 units for power each quarter find the least number of units he/she would need to use for lighting in order to make RATE B the cheaper rate? ANSWER: 286 Please

OpenStudy (perl):

what does 10 c mean here

OpenStudy (anonymous):

Hi ! sorry it means cents. charged 10 cents per unit

OpenStudy (perl):

ok :)

OpenStudy (perl):

Let x = # light units used y = # power units used Cost Rate A = .10x + .02 y Cost Rate B = 10 + .03x + .03y

OpenStudy (texaschic101):

0.10L + .02(1000) < 10 + .03(1000) I must be setting it up wrong....because this does not equal 256..it equals 200

OpenStudy (texaschic101):

oops..not 256...I meant does not equal 286

OpenStudy (texaschic101):

Rate B says .03 for EITHER lighting or power......so I just subbed in power

OpenStudy (texaschic101):

go with what perl said....that equals 286 when rounded

OpenStudy (anonymous):

Thank you !! I'm still going through it ..

OpenStudy (texaschic101):

only I don't think I would put the equal sign in there because you want rate B less then...not less then or equal

OpenStudy (perl):



OpenStudy (perl):

right :) woops

OpenStudy (perl):

Let x = # light units used y = # power units used Cost with Rate A = .10x + .02 y Cost with Rate B = 10 + .03x + .03y We are given y = 1000 Cost Rate A = .10x + .02*1000 Cost Rate B = 10 + .03x + .03*(1000) Cost Rate A = .10x + 20 Cost Rate B = .03x + 40 Solve , when is cost rate B <= cost rate A Solve: .03x + 40 < .10x + 20 the solution is x > 286 , rounding up

OpenStudy (perl):

You can see in the graph that rate A starts out as the cheaper rate, but 286 units of light is the tipping point. For 286 units of light or greater, rate B is cheater.

OpenStudy (perl):

actually 285.714 is the tipping point

OpenStudy (perl):

we can solve that inequality algebraically

OpenStudy (anonymous):

Wow thank you so much!! I can see clearly how you have achieved this result. Thank you very much for responding. I really appreciate it, thank you! I am having trouble with all these types of question ie linear equation word problems. Every time i really struggle to form the equation. Solving the equation itself is the easy part. I spent all day yesterday and so far today dedicated to practicing similar problems online. Thats how i found this site (through 'Pauls Online Maths' webpage - doing his linear equation stuff). Anyway THANK YOU SO SO MUCH :D

OpenStudy (perl):

your welcome :)

OpenStudy (texaschic101):

thats why perl is a qualified helper......very good at explaining stuff

OpenStudy (anonymous):

I wish i could come up with the equation in the first place!

OpenStudy (texaschic101):

thats what I kept messing up on...lol. To me, thats the hardest part

OpenStudy (anonymous):

Do you believe its a practice thing? Because if Im honest I dont think I'm getting any better at it ... Some things are like that, I just want to know an honest answer from someone who knows

OpenStudy (texaschic101):

its practice and paying special attention to the wording of the problem

OpenStudy (texaschic101):

because some word problems are tricky and they throw info in the problem that is not even needed to solve the problem

OpenStudy (anonymous):

heh - i hope so. Its like translating another entire language or something. i just cant see how to make what they're asking in maths terms

OpenStudy (perl):

make sure you understand each step in the solution, thats how i learn for the next problem. each problem you solve gives you tools to work on other problems.

OpenStudy (texaschic101):

thats about right...take it step by step

OpenStudy (xapproachesinfinity):

surely it is the amount of practice and attention to pay to learn them i usually have difficulties with wording problems and that's because of lack of practiced problems i take sometime to understand what i have to do and then attack the problem sometimes they are tricky so once you learn that trick from a problem it gets to next one and so on :) just put a lot of effort

OpenStudy (anonymous):

Perl - that is golden advice, i will try to apply it the next time I dont know where to start heh. (which will no doubt be in about 3minutes - ie after I've written outy the next one to practice on)

OpenStudy (xapproachesinfinity):

believe you will find yourself masted it! just good amount of practice :)

OpenStudy (anonymous):

I hope so xapproachesinfinity, thanks... they do my head in but hopefully soon it will start to turn around :)

OpenStudy (perl):

also try to pick out 'forms' or ideas that recur in problems. like in this problem we have, to get cost, cost per unit * # units cost / unit * # units , notice that the the units cancel

OpenStudy (anonymous):

yeah right. hmmmmmm :D

OpenStudy (anonymous):

Anything else ??? Hahaha :)

OpenStudy (perl):

and the fixed charge of 10, well we just add to add that to B

OpenStudy (perl):

Cost of plan A = (cost per unit light) * (# units of light ) + (cost per unit power) * (# units power )

OpenStudy (anonymous):

yes I see now that too is a hint, to add it. And the next line you're helping me with re the cost i can see too here that you are really breaking it all right down. Thank you for going through it with me

OpenStudy (perl):

Cost of plan B = (cost per unit light) * (# units of light ) + (cost per unit power) * (# units power ) + fixed charge

OpenStudy (perl):

$$ \Large{ \frac {cost }{unit } \cdot \#~ units = \frac {cost }{\cancel{unit} } \cdot \cancel{ \# ~units} \\~\\ } $$

OpenStudy (anonymous):

Im re writing the question out in my workbook and dissecting it/adding all your comments and notes to try and expose the method behind extracting the ACTUAL question and FORMING THE EQUATION

OpenStudy (perl):

there you go :) then next time you will be better prepared to translate the world problem into math